OpenStudy (anonymous):

I need to find the zeros for the function: X^4+X^3-15X^2-16X-16

OpenStudy (anonymous):

caclulator? graph and calculate answers.

OpenStudy (anonymous):

qith rational zero test?

OpenStudy (anonymous):

equate the function to zero an keep factorizing it until you reach 4 factors and those are the zeroes...for factorizing, u can use hit and trial or also you can use quadratic method in which consider u=x^2 and solve...

OpenStudy (anonymous):

I did the synthetic math and figured out that 4 is the number that worked but I don't know what to do after that. This is without calculator.

OpenStudy (anonymous):

@sdgrey1 ya so now......use x-4 as a factor and factorize the polynomial

OpenStudy (anonymous):

-7+8x-12+4x16=-12

OpenStudy (anonymous):

ok so I factored out (x-4) and was left with x^3+5x^2+5x+4. what now

OpenStudy (anonymous):

-7+8x-12+4x16=-12

OpenStudy (anonymous):

me ajuda -7+8x-12+4x16=-12

OpenStudy (anonymous):

@sdgrey1 you have to again use synthetic division to find the other factor and you will get a quadratic equation then you factorize that..........

OpenStudy (anonymous):

ok let me try that

OpenStudy (anonymous):

ok that worked for me. thank you.

OpenStudy (anonymous):

Welcome....